### v1

# Feature Representation Transferring to Lightweight Models via Perception Coherence

Abstract
In this paper, we propose a method for transferring feature representation to lightweight student models from larger teacher models. We mathematically define a new notion called perception coherence. Based on this notion, we propose a loss function, which takes into account the dissimilarities between data points in feature space through their ranking. At a high level, by minimizing this loss function, the student model learns to mimic how the teacher model perceives inputs. More precisely, our method is motivated by the fact that the representational capacity of the student model is weaker than the teacher model. Hence, we aim to develop a new method allowing for a better relaxation. This means that, the student model does not need to preserve the absolute geometry of the teacher one, while preserving global coherence through dissimilarity ranking. Our theoretical insights provide a probabilistic perspective on the process of feature representation transfer. Our experiments results show that our method outperforms or achieves on-par performance compared to strong baseline methods for representation transferring.
1 Introduction
The need of lightweight models and Knowledge distilling (KD). Deep learning models are at the forefront of performance in different domains and tasks, such as classification he2016deep ; tan2019efficientnet ; tan2021efficientnetv2 ; he2020momentum or object detection ren2015faster ; lin2017feature ; lin2017focal ; he2020momentum . However, as the performance of neural networks increases, so does the cost of increasingly large models. In many applications where resources are limited (e.g. mobile devices) or one needs a fast execution, it is preferable to use lightweight models. Among the many approaches to obtain lightweight models, KD buciluǎ2006model ; hinton2015distilling ; wang2021knowledge stands out as a particularly interesting direction. A KD setup essentially consists of a teacher model and a student model. The student model (here the targeted lightweight model) is generally smaller than the teacher model, and therefore more efficient for the task in question. At a high level, the KD process consists in teaching the student model so that it mimics the way the teacher model understands the input. This is usually achieved by using a so-called transfer dataset. The student model is trained to behave like the teacher model on this set in some way . Many previous works have shown that KD improves the generalization capacity of the student model hinton2015distilling ; chen2017learning ; zhu2018knowledge ; beyer2022knowledge ; huang2022knowledge . Different up-to-date KD approaches will be discussed in Section 2.
Feature representation transferring. Classical KD techniques generally focus on matching the outputs of student and teacher models, typically for classification tasks hinton2015distilling . However, this approach has limitations, such as requiring the same number of classes between the teacher and student models. To address these issues, previous works passalis2018learning ; passalis2020probabilistic ; tiancontrastive propose to directly transfer the feature space representations through learning. This enables the student model to better capture the input-input relations represented in the feature space of the teacher model, by preserving its geometry and distribution. This leads to more effective knowledge transfer. Furthermore, note that this approach is an unsupervised transferring method, as it only uses a transfer set without labels. As a consequence, the learned features can be used for various downstream tasks. Because of these appealing properties, our method aligns with this last approach.
Motivation and an overview of our method. Generally, the student model is much smaller than the teacher model. Thus, in terms of representational power, the student model cannot learn to produce features with the exact same geometry as the teacher model. Hence, we aim to develop a method such that (1) the student model does not need to copy the feature representation of the teacher model; (2) the learned representation retains a certain overall consistency with the teacher model’s feature representation and (3) the method works for representation transferring between spaces having different dimensions. For these purposes, we develop a novel fairly simple notion called perception coherence, that is explained subsequently. To begin with, let us consider a reference input $x$ and a set of $n$ point $\{x_{i}\}_{i}^{n}$ (called compared set). These points are fed into the student and teacher models and are mapped into the feature space of each model. The feature space reflects how the models perceive the inputs. Placing ourselves in the two feature spaces, let $d_{1}$ and $d_{2}$ be dissimilarity measures for the teacher and student models, respectively. Then, if $d_{1}(x,x_{i})\leq d_{1}(x,x_{j})$ for some $i,j\in\{1,2,\dotsc,n\}$ ($i\neq j$), we expect that $d_{2}(x,x_{i})\leq d_{2}(x,x_{j})$ also holds. Less formally, if the teacher model perceives $x_{i}$ to be more similar to $x$ than $x_{j}$, we expect the student to have the same perception. We refer to this as the property of perception coherence. By adopting this behavior for different reference points and compared sets, the student model learns to have a coherent perception with respect to the teacher model. The general idea of our method is illustrated in Figure 1.1. Note that we do not measure how much more similar $x_{i}$ is to $x$ than $x_{j}$, but we only rank the dissimilarities. This allows flexibility, as the student does not need to match the teacher’s geometry or distribution exactly. However, the ranking operation is non-differentiable. Hence, we introduce a simple differentiable soft ranking function (Eq. (3.3)). Our contribution can be summarized as follows:
  • We propose a new method for distilling knowledge to lightweight models, based on deep representation transferring. To do so, we develop a novel probabilistic concept called perception coherence. This leads to a novel easy-to-implement loss function (Eq. (3.4)), based on a soft ranking operation.
  • Our theoretical insights via a probabilistic framework provide a new way of understanding representation transfer between models through local and global dissimilarity ranking in the feature spaces between transfer domains. More particularly, we show that a larger global expected coherence level leads to a higher chance that the learned representation respects the dissimilarity ranking of the teacher one.
  • We conduct experiments on retrieval and classification tasks to compare our method with strong baseline methods. The results show that our method outperforms or achieves competitive performance with state-of-the-art methods on lightweight representation transferring settings. Our code is available at anonymized-code.
Figure 1.1: General scheme of our method. In each mini-batch, we use each input as the reference point to compute the dissimilarity to the others (compared set). Then, using loss function in Eq. (3.4), the student model is trained to respect the perception coherence. Notice that $f_{1}$ and $f_{2}$ can be the whole model (when using the output layer) or a part of the model (when using an intermediate layer).
2 Related Work
KD based on soft label produced by teacher model. This approach has been studied since the early days of neural networks buciluǎ2006model . Therein, the authors show that knowledge can be transferred from a large ensemble of models to a single small model. A standard KD method consists in using output of classification models hinton2015distilling ; tang2016recurrent ; huang2022knowledge , with a temperature, leading to smoother soft labels. Based on the same spirit, chan2015transferring proposes to match the output distributions using Kullback-Leibler divergence. Soft label was also used in the context of domain adaption tzeng2015simultaneous . While all these methods show the benefits of soft labels, the main limit of this approach is the need of having the same number of classes between models.
KD based on distance matching. In romero2014fitnets , the authors propose to minimize the distance between student and teacher’s features to guide the student model during training. However, in the case of dimension mismatch between the models, one need to pass through a linear transformation, which leads to substantial information loss. In yim2017gift , the authors propose to minimize distance of the so-call FSP matrix between teacher and student models (capturing inter-layer relations). More advanced methods consist in minimizing the distance between the derivatives w.r.t. to the features czarnecki2017sobolev or between the attention map of student and teacher zagoruyko2016paying . All these approaches require the dimension match, and need some form of interpolation in the case of dimension mismatch. Overcoming the drawback of dimension mismatch, passalis2018unsupervised proposes to match the the pair-wise similarity matrix of the training points (in the transfer set) between the student and teacher models. The method proposed in yu2019learning is to match the distance between pairs of samples. However, this brute force approach is not likely to work as the student model is much smaller than the teacher one (as discussed in passalis2020probabilistic ).
Feature representation transfer. Some previous works treat the representation transfer as the problem of distribution matching, based on Maximum Mean Discrepancy huang2017like or adversarial training belagiannis2018adversarial . The latter method is inspired by Generative Adversarial Networks (GANs) goodfellow2014generative . One major drawback of distribution matching is the need of having the same dimension between the two feature spaces. Probabilistic frameworks based on mutual information are proposed in ahn2019variational ; tiancontrastive . However, the method in ahn2019variational requires a Gaussian assumption (this is not necessarily satisfied in practice) while the method tiancontrastive requires the input features to have the same dimension (hence requiring a linear transformation, which may reduce the useful information). Overcoming the requirement of dimension match, park2019relational proposes to transfer the input relations using distance-wise and angle-wise distillation loss. However, the structural information is quite limited, as it only considers a single triple of points at a time. Overcoming this drawback, passalis2018learning proposes a method of probabilistic knowledge transfer (PKT). The high level idea of this last method is to transfer the local geometry learned by the teacher model based on kernel techniques. Based on this work, passalis2020heterogeneous ; passalis2020probabilistic basically share the same idea, but they apply the technique for multiples intermediate layers with an auxiliary model. The benefits of using auxiliary models are discussed in mirzadeh2020improved . As in passalis2018learning ; passalis2020heterogeneous ; passalis2020probabilistic , our method takes into account the relation between the inputs and allows representation transfer between space having different dimensions. However, in passalis2018learning ; passalis2020heterogeneous ; passalis2020probabilistic , a common kernel is used for all points (e.g. a Gaussian one with a fixed bandwidth), which cannot adapt to local density variation. This contrasts with our method, where we only rank the dissimilarities but do not use the dissimilarity magnitude. Thus, our method overcomes the problem of local scale variations, allowing itself to adapt flexibly to the local density.
3 Our method
3.1 Perception coherence and other notions
We consider a task where the input space is denoted by $\mathcal{X}$ associated with the underlying distribution $\mathcal{D}_{\mathcal{X}}$. Let us consider two feature extractors $f_{1}:\mathcal{X}\mapsto\mathcal{F}_{1}$ and $f_{2}:\mathcal{X}\mapsto\mathcal{F}_{2}$, endowed with symmetric dissimilarity metrics $d_{f_{1}}:\mathcal{F}_{1}\times\mathcal{F}_{1}\mapsto\mathbb{R}$ and $d_{f_{2}}:\mathcal{F}_{2}\times\mathcal{F}_{2}\mapsto\mathbb{R}$, respectively. For ease of following, we can think of $f_{1}$ and $f_{2}$ as teacher and student models, respectively. For brevity, for any $x,x^{\prime}\in\mathcal{X}$, let $d_{1}(x,x^{\prime})=d_{f_{1}}(f_{1}(x),f_{1}(x^{\prime}))$ and $d_{2}(x,x^{\prime})=d_{f_{2}}(f_{2}(x),f_{2}(x^{\prime}))$.
At a high level, if $f_{1}$ perceives $x_{1}$ to be more similar to $x$ than $x_{2}$, then it should be also the case for $f_{2}$ after the knowledge distilling. This is represented through the feature space of each model, endowed with their corresponding dissimilarity metric. We present this intuition via a more formal concept named absolute perception coherence as follows.
Definition 1 (Absolute perception coherence).
We say that $f_{2}$ is absolutely perception coherent with $f_{1}$ at $x\in\mathcal{X}$ if $d_{1}(x,x_{2})\geq d_{1}(x,x_{1}))\Leftrightarrow d_{2}(x,x_{2})\geq d_{2}(x,x_{1})$, for any $x_{1},x_{2}\in\mathcal{X}$.
However, the absolute perception coherence is too strict. Thus, we shall introduce a more relaxed notion subsequently, based on a probabilistic framework by using the distribution $\mathcal{D}_{\mathcal{X}}$. To begin with, given $x,x^{\prime}\in\mathcal{X}$, let us define the cumulative functions as follows
$F_{1}(x,x^{\prime}):=\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(x,x^{\prime})\right),\;F_{2}(x,x^{\prime}):=\mathbb{P}_{X}(d_{2}(x,X)\leq d_{2}(x,x^{\prime}))\ .$ (3.1)
Remark 1.
We have the following remarks on the relation between the previous function and the ordering of dissimilarities to a given reference point $x$.
  • $F_{i}(x,x_{1})\leq F_{i}(x,x_{2})\Leftrightarrow d_{i}(x,x_{1})\leq d_{i}(x,x_{2}),\;\text{for $i\in\{1,2\}$}$.
  • A small value of $F_{i}(x,x^{\prime})$ means that $x^{\prime}$ is probabilistically close to $x$. In other words, if we choose randomly another point $\widetilde{x}$ according to $\mathcal{D}_{\mathcal{X}}$, there is a high chance that $d_{i}(x,\widetilde{x})\geq d_{i}(x,x^{\prime})$.
  • Similarly, a large value of $F_{i}(x,x^{\prime})$ means that $x^{\prime}$ is probabilistically far from $x$.
  • Thus, for each model $f_{i}$, $F_{i}(x,x^{\prime})$ represents “probabilistic distance” between $x$ and $x^{\prime}$.
Remark 2 (Relation between cumulative function and ranking).
Let us consider a discrete distribution composed of $N+1$ examples $\{x\}\bigcup\{x_{i}\}_{i=1}^{N}$ and the extractor $f$ endowed with a dissimilarity metric $d$. We consider $x$ as the reference point and we compute dissimilarity between $x$ and other points $x_{i}$’s to obtain $D:=\{d(x,x_{i})\}_{i=1}^{N}$. Assuming there is no ties, i.e., $d(x,x_{i})\neq d(x,x_{j})$ for all $i\neq j$, then, the cumulative function $F(x,x_{i})$ is the normalized rank of $d(x,x_{i})$ in the set $D$.
Using the cumulative function of Eq. (3.1), we define the perception coherence level as follows.
Definition 2 (Perception coherence level).
The perception coherence level of $f_{2}$ with respect to (w.r.t.) $f_{1}$ at $x\in\mathcal{X}$ is defined as
$\phi_{f_{1},f_{2}}(x):=1-\mathbb{E}_{X}\left[\left|F_{1}(x,X)-F_{2}(x,X)\right|\right].$ (3.2)
Remark 3.
The coherence level is always valued in $[0,1]$. Indeed, $0\leq\left|F_{1}(x,x^{\prime})-F_{2}(x,x^{\prime})\right|\leq 1$ for all $x,x^{\prime}\in\mathcal{X}$. Hence, $0\leq\mathbb{E}_{X}\left[\left|F_{1}(x,X)-F_{2}(x,X)\right|\right]\leq 1$.
Proposition 1.
If $f_{2}$ is absolutely perception coherent with $f_{1}$ at $x$, then $\phi_{f_{1},f_{2}}(x)=1$.
The proof in this proposition can be found in Appendix A. Proposition 1 tells that if we have an absolute perception coherence between the two models at a point $x$, the coherence level is equal to 1 at $x$. Now, a natural question to ask is whether the reverse statement holds. In short, our theoretical analysis yields a positive answer to this question. This will be presented in Section 3.3.
The coherence level naturally gives rise to the following definitions for the local and global perception coherence.
Definition 3 (Local $\alpha$-perception coherence).
We say that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ if $\phi_{f_{1},f_{2}}(x)\geq\alpha$.
Definition 4 (Global $\alpha$-perception coherence).
We say that $f_{2}$ is globally $\alpha$-perception coherent with $f_{1}$ if $\mathbb{E}_{X}[\phi_{f_{1},f_{2}}(X)]\geq\alpha$. That is, the global expected coherence level is at least $\alpha$.
3.2 Practical implementation with a differentiable ranking operation
To implement our method, we use a simple Monte-Carlo sampling, processing by mini-batch, like any standard machine learning algorithm. Consider a training batch :=$\mathcal{B}:=\{x_{i}\}_{i=1}^{B}$. For each $f\in\{f_{1},f_{2}\}$ endowed with the metrics $d\in\{d_{f_{1}},d_{f_{2}}\}$, we first compute the pair-wise dissimilarity $d(x_{i},x_{j})$ for all $i\neq j$. Let $d_{ij}$ denote $d(x_{i},x_{j})$ for brevity. Let us denote $D^{f}_{i}(\mathcal{B}):=(d_{ij})_{j=1}^{B}$. Let $\mathcal{R}^{f}_{i}(\mathcal{B})$ be the ranking of $D^{f}_{i}(\mathcal{B})$. That is, the $j^{th}$ component of $\mathcal{R}^{f}_{i}(\mathcal{B})$ is the rank of $d_{ij}$ in $D^{f}_{i}(\mathcal{B})$. Notice that $\mathcal{R}^{f}_{i}(\mathcal{B})$ is a permutation of $\{1,2\cdots,B\}$ (so, $\mathcal{R}^{f}_{i}(\mathcal{B})\in\mathbb{R}^{B}$). Based on Remark 2, intuitively, with sufficiently large $B$ [^1] Through our qualitative observation, for $B$ from $\sim 10$, results start to stabilize. See Appendix E.4 for results and discussions., $\frac{\mathcal{R}_{i}^{f_{1}}(\mathcal{B})}{B}\approx\left(F_{1}(x_{i},x_{j})\right)_{j=1}^{B}$ (similarly for $f_{2}$). Hence, to empirically maximize the perception coherence level (Eq. (3.2)), we can minimize
$\mathcal{L}(f_{1},f_{2};\mathcal{B}):=\frac{1}{B}\sum_{i=1}^{B}\left\|\frac{\mathcal{R}_{i}^{f_{1}}(\mathcal{B})}{B}-\frac{\mathcal{R}_{i}^{f_{2}}(\mathcal{B})}{B}\right\|^{2}=\frac{1}{B^{3}}\sum_{i=1}^{B}\left\|\mathcal{R}_{i}^{f_{1}}(\mathcal{B})-\mathcal{R}_{i}^{f_{2}}(\mathcal{B})\right\|^{2}\ .$
We use here the squared Euclidean norm for its simplicity, but other metrics can be used to compute the difference between $\mathcal{R}_{i}^{f_{1}}(\mathcal{B})$ and $\mathcal{R}_{i}^{f_{2}}(\mathcal{B})$. Besides, notice that each $i\in\{1,2\cdots,B\}$ corresponds to $x_{i}$ being the reference point. Hence, averaging over $i$ in the batch allows for maximizing the global expected coherence level (the expectation is taken over different reference points, see Definition 4 for recalling). However, the ranking operation is not differentiable. Hence, we propose a simple soft ranking operation based on sigmoid function. Assuming that there is no ties, i.e., $d_{ij}\neq d_{ik}$, for all $j\neq k$, that generally holds in practice. In this case, we notice that the rank of $d_{ij}$ in $D_{i}(\mathcal{B})$ can be computed as $r(d_{ij})=\sum_{k=1}^{B}\mathds{1}_{\{d_{ik}\leq d_{ij}\}}$. Hence, we can soften the ranking operation as follows
$\widetilde{r}(d_{ij})=\sum_{k=1}^{B}\Lambda\left(\frac{d_{ij}-d_{ik}}{\tau}\right)\ ,$ (3.3)
where $\Lambda$ is the sigmoid function and $\tau<1$ is the temperature playing the role of scaling. The above function approaches the step function as $\tau$ tends to 00. Hence, this allows us to conveniently approximate the ranking operation. Let ~$\widetilde{\mathcal{R}}_{i}(\mathcal{B})$ be the corresponding ranking obtained with this softened version. So, our loss function can be written as,
$\mathcal{L}_{\text{ours}}(f_{1},f_{2};\mathcal{B}):=\frac{1}{B^{3}}\sum_{i=1}^{B}\left\|\widetilde{\mathcal{R}}_{i}^{f_{1}}(\mathcal{B})-\widetilde{\mathcal{R}}_{i}^{f_{2}}(\mathcal{B})\right\|^{2}\ .$ (3.4)
3.3 Theoretical insights
Throughout our theoretical analysis, we make the following assumption.
Assumption 1.
Let $X$ and $X^{\prime}$ be the i.i.d random variables following the law $\mathcal{D}_{\mathcal{X}}$. We assume that for any $x\in\mathcal{X}$, $\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|=\left|F_{2}(x,X)-F_{1}(x,X)\right|\right)=0$. We remark that this assumption is reasonable in the context of continuous variable modeling.
In this section, we first derive theoretical insights about the local perception coherence at a given point $x$ in Section 3.3.1. Next, we also investigate about the global coherence in Section 3.3.2. Finally, we study the stability of the perception coherence around a given local region in Section 3.3.3.
3.3.1 Local perception coherence
Theorems 3.1 and 3.2 provide insight into how well the student model $f_{2}$ mimics the teacher model $f_{1}$, given the local perception coherence level.
Theorem 3.1 (Local rank preservation).
Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ (i.e. $\phi_{f_{1},f_{2}}(x)\geq\alpha$). Then, for any $\varepsilon_{2}>\varepsilon_{1}>0$, we have
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (3.5)
where $C_{\varepsilon_{1}}$ is a positive constant depending on $\varepsilon_{1}$.
The proof is in Appendix B.1. Consider the case where $\alpha$ approaches 1, then the right-hand side (RHS) of Eq. (3.5) approaches 0. Theorem 3.5 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives them to have close dissimilarities w.r.t. a given point $x$ (i.e., their dissimilarities differ by at most $\varepsilon_{1}$), then there is a high chance that it is also the case for the student model $f_{2}$. That is, $f_{2}$ also perceives that their dissimilarities differ by at most $\varepsilon_{1}$. This is because if $\alpha=1$, the RHS is equal to 00 for any $\varepsilon_{2}>\varepsilon_{1}$. While this theorem tells us only about the difference between the dissimilarities (absolute value), Theorem 3.2 also considers the ordering of the dissimilarities.
Theorem 3.2 (Local relative order preservation).
Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ ($\phi_{f_{1},f_{2}}(x)\geq\alpha$). Then, for any $\varepsilon>0$, we have
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (3.6)
where $C$ is a positive constant.
The proof of this theorem can be found in Appendix B.2. Notice that the RHS of Eq. (3.6) approaches 00 as $\alpha$ approaches $1$ for any $\varepsilon>0$. Hence, in this case, Theorem 3.2 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives $x_{1}$ to be more similar to $x$ than $x_{2}$ (i.e., $F_{1}(x,x_{1})\leq F_{1}(x,x_{2})$), then probability that $f_{2}$ perceives $x_{1}$ to be less similar to $x$ than $x_{2}$ is upper bounded. In the case where $\alpha=1$, this probability is 00, that can be formally stated as follows.
Corollary 3.1.
Assume that $f_{2}$ is $1$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ ($\phi_{f_{1},f_{2}}(x)=1$). Then, for any $\varepsilon>0$, we have
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)=0\ .$
Note that this holds for any $\varepsilon>0$. Thus, by letting $\varepsilon\to 0$, the above corollary tells that, the probability that $f_{2}$ perceives $x_{1}$ to be less similar to $x$ than $x_{2}$ is null. That is, the student model $f_{2}$ learns to have the same perception as the teacher model $f_{1}$.
3.3.2 Global perception coherence
In the above section, we analyzed the perception coherence at a reference point $x$, knowing that the local coherence level at $x$ is at least $\alpha$. Now, a natural question concerning the global coherence follows: What conclusions can be drawn from the global expected coherence level when the reference point $x$ is not fixed? Theorem 3.3 formally answers this question.
Theorem 3.3.
Assume that $f_{2}$ is globally $\alpha$-perception coherent with $f_{1}$ (i.e., $\mathbb{E}_{X}[\phi_{f_{1},f_{2}}(X)]\geq\alpha$). Let $X_{1}$, $X_{2}$ and $X_{3}$ be i.i.d. random variables following the law $\mathcal{D}_{\mathcal{X}}$. Then,
  1. 1.
    Global rank preservation. For any $\varepsilon_{2}>\varepsilon_{1}>0$, we have
    $\mathbb{P}\left(\left|F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\right|\geq\varepsilon_{2}\Bigm{|}\left|F_{1}(X_{1},X_{2})-F_{1}(X_{1},X_{3})\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$
    where $C_{\varepsilon_{1}}$ is a positive constant depending only on $\varepsilon_{1}$.
  2. 2.
    Global relative order preservation. For any $\varepsilon>0$, we have
    $\mathbb{P}\left(F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\geq\varepsilon\Bigm{|}F_{1}(X_{1},X_{2})\leq F_{1}(X_{1},X_{3})\right)\leq\frac{C(1-\alpha)}{\varepsilon}\ ,$
    where $C$ is a positive constant.
The proof of this theorem can be found in Appendix C.1. Theorem 3.3 shows that, even without fixing the reference point, knowing the global expected coherence guarantees a bound. This result can be interpreted similarly to the local perception case in Theorems 3.1 and 3.2.
3.3.3 Stability of perception coherence around a local region
We now study the fluctuation around a local region $\mathcal{A}\subseteq\mathcal{X}$. We seek to answer the following question: If the local perception coherence level is at least $\alpha$ in the region $\mathcal{A}$, how stable is the coherence level under perturbations around $\mathcal{A}$? In short, our theoretical result in Theorem 3.4 will show that the fluctuation of the coherence is bounded by the perturbation level around $\mathcal{A}$. For this, we make the following assumptions.
Assumption 2.
We make the following assumptions.
  • The dissimilarities of both models are symmetric and satisfy the triangular inequality.
  • For $\delta\geq 0$, let $\rho(\delta)=\max_{i\in\{1,2\}}\sup_{x\in\mathcal{X},d\in\mathbb{R}}\mathbb{P}_{X}\left(d_{i}(x,X)\in[d,d+\delta]\right)$. We assume that $\rho(\delta)=O(\delta)$ as $\delta\to 0$.
  • For $i\in\{1,2\}$, $x,x^{\prime}\in\mathcal{X}$, $d_{i}(x,x^{\prime})=O(F_{i}(x,x^{\prime}))$ as $F_{i}(x,x^{\prime})\to 0$.
Let $F(x^{\prime},x)=\max(F_{1}(x^{\prime},x),F_{2}(x^{\prime},x))$. Intuitively, a small value of $F(x^{\prime},x)$ signifies that $x$ and $x^{\prime}$ are probabilistically close to each other in the feature spaces of both models $f_{1}$ and $f_{2}$. Then, the relation between the fluctuation of the perception coherence and the perturbation around $\mathcal{A}$ is stated as follows.
Theorem 3.4.
Consider a non-empty set $\mathcal{A}\subseteq\mathcal{X}$ such that $\phi_{f_{1},f_{2}}(x)\geq\alpha,\;\forall x\in\mathcal{A}$. For $0<\delta\leq 1$, let $\mathcal{A}_{\delta}=\{x\in\mathcal{X}:\min_{x^{\prime}\in\mathcal{A}}F(x^{\prime},x)\leq\delta\}$. Then, for any $0<\varepsilon\leq\alpha$, we have
$\mathbb{P}_{X}\left(\phi_{f_{1},f_{2}}(X)\leq\alpha-\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)\leq\frac{O(\delta)}{\varepsilon C_{\mathcal{A}}}\ ,$ (3.7)
where $C_{\mathcal{A}}$ is a positive constant depending on $\mathcal{A}$.
The proof can be found in Appendix D. Notice that $\mathcal{A}_{\delta}$ is the perturbed set around $\mathcal{A}$ at perturbation level $\delta$ ($\mathcal{A}\subseteq\mathcal{A}_{\delta}$). Theorem 3.4 states that if the coherence level in $\mathcal{A}$ is at least $\alpha$, then the probability that the coherence level of a point around $\mathcal{A}$ drops below $\alpha-\varepsilon$ is bounded above by $O(\delta)$. This holds for any $\varepsilon>0$, indicating that the coherence level does not decrease significantly as we expand the region $\mathcal{A}_{\delta}$ (by increasing $\delta$). Thus, as the perturbation $\delta\to 0$, the upper bound also tends to zero, characterized by $O(\delta)$. This ensures the stability of coherence level within each local region.
4 Experiments
In this section, we first conduct experiments on $2D$ and $3D$ data for proof-of-concept (Section 4.1). Then, in Section 4.2, we conduct experiments on lightweight settings to evaluate the quality of the learned representation through retrieval tasks. Finally, we demonstrate how our method can boost the performance of student models in classification task in Section 4.3. Without further mention, we use cosine dissimilarity as the dissimilarity metric on neural networks (defined in Appendix E). For brevity, the details for all experiments can also be found in Appendix E.
4.1 Proof-of-concept with experiments on $\mathbf{2D}$ and $\mathbf{3D}$ toy examples
In this section, we conduct experiments on $2D$ and $3D$ data, without neural networks, to qualitatively demonstrate how our method works. For this, we use a reference data configuration $\{x^{1}_{i}\}_{i=1}^{N}$ (can be seen as the teacher configuration, $N$ is the number of data points). Then, we randomly initialize another configuration $\{x^{2}_{i}\}_{i=1}^{N}$ (student configuration), where each point $x^{2}_{i}$ is associated with a fixed point $x^{1}_{i}$. More formally, for each $i\in\{1,2\cdots N\}$, $f_{1}(i)=x^{1}_{i}$ (fixed) and $f_{2}(i)=x^{2}_{i}$. We then apply our loss function from Eq. (3.4), where the dissimilarity metric is the Euclidean distance. All experiment details are in Appendix E.1. The evolution of the student configurations in different setups are shown in Fig. 4.1 and 4.2. We intentionally set the teacher and student configurations in different scales of magnitude to show the effectiveness of the method (teacher configuration has data points in scale $\sim 1$ whereas student configuration is in scale $\sim 10$). We see that our method results in a very smooth configuration transfer, despite the scale difference. From Fig. 4.1, we remark that, while the exact geometry is not completely preserved, the learned configuration preserves a global structural coherence. The same phenomenon is also observed in the case where teacher and student configurations have different dimensions ($3D\to 2D$) in Fig. 4.2. Another experiment with more clusters is in Appendix E.1. Thus, this qualitatively proves the effectiveness of our method.
Figure 4.1: Transferring process on $2D$ datasets. Top row represents the teacher configuration (fixed), bottom row represents the evolution of the student configuration along the transfer process (from left to right). Each gray line associates $x^{2}_{i}$ with $x^{1}_{i}$ (for the same $i$). We observe that the learned configuration preserves a global structural coherence without preserving completely the geometry.
(a) $3D$ reference configuration.
(b) Evolution along the transferring process of learned configuration in $2D$.
Figure 4.2: Transferring process from $3D$ to $2D$ data.
4.2 Learning Metric Evaluation on Neural Networks
In this section, we perform experiments on CIFAR10 cifar10 and the subset of the first 30 classes of CUB-200 wah2011caltech with very small student models. Following the previous works (e.g. passalis2018learning ; passalis2020probabilistic ; passalis2020heterogeneous ), we evaluate the quality of the learned representation using the content-based image retrieval setup. We used the official code released with the work passalis2020heterogeneous to evaluate different metrics, including the interpolated mean Average Precision (mAP) (at the standard 11-recall points) and the top-$k$ precision. For a fair comparison, we use the same teacher and student model architecture as in passalis2020heterogeneous . The student models are very small, composed of only 3 convolutional layers and a fully connected layer. All details are in Appendix E.2. To conduct the experiments, for each dataset CIFAR10 and CUB-200, we first train the teacher model on the training set. Then, the transfer set is only the training set without label. Then, the retrieval setup is used to evaluate the quality of the learned representation. The database contains the training set and the test set is used for evaluating the performance of each model. If the student is trained to learn useful features, then the test examples should be close to the training examples (database) of the same class. Hence, a higher retrieval score indicates a better quality of the learned representation. The results for CIFAR10 and CUB-200 are shown in Tables 4.1 and 4.2, respectively. Our method is also compared with classical methods, including the standard knowledge distilling (KD) hinton2015distilling , FitNet romero2014fitnets , metric knowledge transfer (MKT) yu2019learning , probabilistic knowledge transfer (PKT) passalis2018learning and Heterogeneous Knowledge Distillation (HKD) passalis2020heterogeneous . The results displayed for all these methods in Tables 4.1 and 4.2 are extracted from passalis2020heterogeneous . Therein, one used the same teacher and student models. For FitNet, MKT, PKT, there are two versions: the first one applies only on a single penultimate layer, while the second one applies on multiple layers (denoted by -H).
Table 4.1: Learning Metric Evaluation on CIFAR10.
Multi-layer or single layer distilling Method Auxiliary Model mAP top-100
Teacher Model (ResNet-18) 91.94 93.50
Single Layer Distilling Ours No 54.25 65.00
KD hinton2015distilling No 40.53 58.56
FitNet romero2014fitnets No 48.99 62.42
MKT yu2019learning No 38.20 52.72
PKT passalis2018learning No 51.56 62.50
Multi-layer Distilling FitNet-H No 46.46 60.59
MKT-H No 43.99 57.63
PKT-H No 51.73 63.01
HKD passalis2020heterogeneous Yes 53.06 64.24
Table 4.2: Learning Metric Evaluation on a subset of the first 30 classes of CUB200.
Multi-layer or single layer distilling Method Auxiliary Model mAP top-10
Teacher Model (ResNet-18) 67.48 74.38
Single Layer Distilling Ours No 28.42 36.55
KD hinton2015distilling No 18.55 26.57
FitNet romero2014fitnets No 15.98 23.41
MKT yu2019learning No 13.39 20.59
PKT passalis2018learning No 18.57 26.70
Multi-layer Distilling FitNet-H No 15.37 22.61
MKT-H No 15.39 22.76
PKT-H No 17.77 25.39
HKD passalis2020heterogeneous Yes 19.01 27.67
From Tables 4.1 and 4.2, we observe that MKT method provides the worst results. This method is composed of 2 main ingredients: (1) minimizing the distance between the student and teacher features and (2) enforcing the student model to result in similar distances between the data points as the teacher model (in the feature space). Hence, in our setting of very small student models, this brute force approach provides the worst results. Our method also outperforms by a large gap the standard approaches of KD (using the soft label) or FitNet (using the distance between the student and teacher features). More advanced methods such as PKT (based on a kernel method) performs better, but it is still outperformed by our method. Besides, HKD is a more advanced version of PKT, where one uses a particular mechanism to transfer the features. However, our method still provides better results. Notice that our method is applied only on the penultimate layer, but outperforms the other methods using multiple layers. Especially on CUB-200, our method outperforms the others by a wide margin. This proves the effectiveness of our method for feature representation transferring.
4.3 Our method as a guide for knowledge distilling: comparing with other methods of feature representation transfer in the context of classification
In the above section, we show how our method helps the student model to produce useful features. In this section, we demonstrate that by transferring the feature representation, our method boosts significantly the performance of the student model in the context of the classification task on CIFAR100 krizhevsky2009learning . While different methods have been proposed and tailored specifically for the classification, we focus mainly on state-of-the-art feature representation transferring methods. Notice that different methods (such as FitNet romero2014fitnets ) apply the transferring on different intermediate layers. However, we propose to apply our method only on the penultimate layer and on the logit one (before applying the softmax function). We also fix all the hyperparameters for training, without further tuning to show the simplicity of our method. All details can be found in Appendix E.3. While different methods have been proposed to effectively transfer knowledge between model of the same architecture type (e.g. ResNet50 $\to$ ResNet18), we intentionally choose pairs of models having different architecture types. More precisely, we use 3 pairs of teacher and student models: ResNet-50 $\to$ MobileNetV2 and ResNet-32x4 $\to$ ShuffleNetV1 / ShuffleNetV2. We follow the same setup and use the same pretrained teacher models officially released with the work tiancontrastive for fair comparisons.
[…]
Method ResNet-50 $\rightarrow$ MobileNetV2 ResNet-32x4 $\rightarrow$ ShuffleNetV1 ResNet-32x4 $\rightarrow$ ShuffleNetV2
Teacher 79.34 79.42 79.42
Student 64.6 70.5 71.82
KD hinton2015distilling 67.35$\pm$0.32 74.07$\pm$0.19 74.45$\pm$0.27
FitNet romero2014fitnets 63.16$\pm$0.47 73.59$\pm$0.15 73.54$\pm$0.22
VID ahn2019variational 67.57$\pm$0.28 73.38$\pm$0.09 73.40$\pm$0.17
RKD park2019relational 64.43$\pm$0.42 72.28$\pm$0.39 73.21$\pm$0.28
PKT passalis2018learning 66.52$\pm$0.33 74.10$\pm$0.25 74.69$\pm$0.34
CRD tiancontrastive 69.11$\pm$0.36 75.11$\pm$0.32 75.65$\pm$0.10
Ours 68.64$\pm$0.10 75.62$\pm$0.28 76.04$\pm$0.25
From Table 4.3, we see that our method boosts significantly the performance of the student models (when not using any KD technique). It also provides much better results than standard techniques such as KD or FitNet. This proves its effectiveness for extracting knowledge from the teacher model. Moreover, it outperforms all the other methods on ResNet-32x4/ShuffleNetV1 and ResNet-32x4 /ShuffleNetV2 benchmarks. On the experiment ResNet-50/ MobileNetV2, our method is slightly outperformed by the method CRD, but outperforms all the other methods. Besides, our method systematically outperforms the PKT method based on kernel techniques (see Section 2). This reinforces our intuition that our method is more effective than applying the same kernel for all points, as discussed in Section 2. Other methods based on mutual information such as Variational Information Distillation (VID) is also outperformed by ours. The method of Contrastive Representation Distillation (CRD) is also based on an information-theoretic framework. Our method provides better or on-par performance compared to CRD. All these remarks prove the effectiveness of our method for KD.
Finally, notice that some different knowledge distilling methods are tailored specifically for classification (e.g. huang2022knowledge ). However, our main objective is to study methods for feature representation transferring, as this class of methods is more generic and applicable on a wider range of tasks (e.g. retrieval). Through our experiments, we aim to show that a good feature representation transfer allows us to effectively distill the knowledge from a neural network.
5 Discussion and Conclusion
Preserving dissimilarity ranks is a flexible requirement, as it does not enforce a full copy of the geometry. However, a question remains theoretically unresolved in this work: How well can a given student model preserve these rankings (as a function of its capacity)? This question remains open in the present work and will be explored in future theoretical works. Besides, notice that by considering the dissimilarity ranking, our method provides structural information of the data manifold. This is related to topological data analysis (TDA), which aims to understand the global data structure. The intersection of TDA and Deep Learning is referred to as topological deep learning (TDL) zia2023topological . Topology studies invariant properties under continuous deformation. In our case, as long as the ranking of relatives distances remains unchanged, we also have invariance under local and global deformation. This allows a better flexibility and stability. One main difference of our method is that, topology studies a configuration based on deterministic approaches, whereas our method proposes a more relaxed view by introducing a probabilistic framework. Besides, by integrating neural networks into this framework, our work is the first attempt to develop TDL for representation transfer and KD.
To sum up, we proposed an original method for deep representation transferring between models having arbitrary feature dimensions. This is achieved based on a novel concept named perception coherence. We propose a probabilistic framework to better understand the proposed notion. The theoretical results show that in the ideal scenario, the student model perceives the inputs in the same way as the teacher model. This provides a new way to theoretically understand the KD process. The experiments show promising results for lightweight KD. Moreover, our method is built generically for any dissimilarity metric for each of the student and teacher models. Hence, depending on application, it would be interesting to investigate task-specific metrics for a best KD, for example, in the multi-domain or multi-modal scenarios. We hope our work paves the way for future advancements in KD, especially based on topological deep learning frameworks.
Broader impact statement
Our work focuses on developing methods to distill knowledge from a neural networks to smaller, more efficient ones. By reducing the computational and memory requirements while maintaining good performance, these techniques have the potential to democratize access to advanced AI systems, making them more accessible across various applications. This could lead to advancements in areas such as on-device AI, or quick personalized services, through improved efficiency and effectiveness of AI technologies.
Appendix A Proof of Proposition 1
Proof.
By definition, we have that $F_{1}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)$ and $F_{2}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$. As $f_{2}$ is absolutely perception coherent with $f_{1}$ at $x$, we obtain that $d_{1}(x,X^{\prime})\leq d_{1}(x,X)\Leftrightarrow d_{2}(x,X^{\prime})\leq d_{2}(x,X)$. Thus, $\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$, which leads to the result. ∎
Appendix B Proof of Theorems 3.1 and 3.2
B.1 Proof of Theorem 3.1
Theorem B.1 (Theorem 3.1 in the main text).
Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ ($\phi_{f_{1},f_{2}}(x)\geq\alpha$). Then, for all $\varepsilon_{2}>\varepsilon_{1}>0$, we have
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (B.1)
where $C_{\varepsilon_{1}}$ is a positive constant depending on $\varepsilon_{1}$.
Proof.
For the sake of brevity, let $A(X,X^{\prime})$ be the event $\left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}$. We have
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ A(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})+F_{1}(x,X^{\prime})-F_{1}(x,X)+F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon_{2}\Bigm{|}A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|+\left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|+\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon_{2}\Bigm{|}A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|+\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon_{2}-\varepsilon_{1}\ \Bigm{|}\ A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\frac{\varepsilon_{2}-\varepsilon_{1}}{2}\ \Bigm{|}\ A(X,X^{\prime})\right)\ .$
Using Lemma 1, we have
$\mathbb{P}_{X,X^{\prime}}\left(\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\frac{\varepsilon_{2}-\varepsilon_{1}}{2}\ \Bigm{|}\ A(X,X^{\prime})\right)$
$\leq\frac{4(1-\alpha)}{(\varepsilon_{2}-\varepsilon_{1})\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}=\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$
where $C_{\varepsilon_{1}}=\frac{4}{\mathbb{P}_{X,X^{\prime}}\left(\left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)}\ .$ Therefore,
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ A(X,X^{\prime})\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ .$
This completes the proof. ∎
B.2 Proof of Theorem 3.2
Theorem B.2 (Theorem 3.2 in the main text).
Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$ ($\phi_{f_{1},f_{2}}(x)\geq\alpha$). Then, for all $\varepsilon>0$, we have
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (B.2)
where $C$ is a positive constant.
Proof.
For the sake of brevity, let $A(X,X^{\prime})$ be the event $F_{1}(x,X^{\prime})\leq F_{1}(x,X)$. We have that
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ A(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)+\left(F_{1}(x,X^{\prime})-F_{1}(x,X)\right)-\left(F_{1}(x,X^{\prime})-F_{1}(x,X)\right)\geq\varepsilon\ \big{|}\ A(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})+F_{1}(x,X)-F_{2}(x,X)+F_{1}(x,X^{\prime})-F_{1}(x,X)\geq\varepsilon\ \big{|}\ A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})+F_{1}(x,X)-F_{2}(x,X)\geq\varepsilon\ \big{|}\ A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(\max\left(F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime}),F_{1}(x,X)-F_{2}(x,X)\right)\geq\varepsilon/2\ \big{|}\ A(X,X^{\prime})\right)$
$\leq\mathbb{P}_{X,X^{\prime}}\left(\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\varepsilon/2\ \big{|}\ A(X,X^{\prime})\right)\ .$
Using Lemma 1, we have
$\mathbb{P}_{X,X^{\prime}}\left(\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\varepsilon/2\ \big{|}\ A(X,X^{\prime})\right)$
$\leq\frac{4(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}=C(1-\alpha)/\varepsilon\ .$
This leads to the result of Theorem 3.2, which completes the proof. ∎
B.3 Lemma 1
Lemma 1.
Let $X$ and $X^{\prime}$ be the i.i.d random variables following the law $\mathcal{D}_{\mathcal{X}}$ and $A(X,X^{\prime})$ be an event depending on these two variables, $A$ is symmetric in $X$ and $X^{\prime}$. Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ at $x\in\mathcal{X}$. Then, for any $\varepsilon>0$, we have that
$\mathbb{P}_{X,X^{\prime}}\left(\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\varepsilon\ \big{|}\ A(X,X^{\prime})\right)$
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}\ .$ (B.3)
Proof.
Let $B(X,X^{\prime})$ be the event $\left|F_{2}(x,X)-F_{1}(x,X)\right|\geq\left|F_{1}(x,X^{\prime})-F_{2}(x,X^{\prime})\right|$. We note that $B(X^{\prime},X)$ is the event $\left|F_{2}(x,X)-F_{1}(x,X)\right|\leq\left|F_{1}(x,X^{\prime})-F_{2}(x,X^{\prime})\right|$. Furthermore, let $C(X,X^{\prime})$ be the event $\max\left(\left|F_{2}(x,X^{\prime})-F_{1}(x,X^{\prime})\right|,\left|F_{1}(x,X)-F_{2}(x,X)\right|\right)\geq\varepsilon$.
$\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap B(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)+\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap\overline{B(X,X^{\prime})}\ \Bigm{|}\ A(X,X^{\prime})\right)\ .$
Using Assumption 1, we have
$\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap\overline{B(X,X^{\prime})}\ \Bigm{|}\ A(X,X^{\prime})\right)=\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap B(X^{\prime},X)\ \Bigm{|}\ A(X,X^{\prime})\right)\ .$
Therefore,
$\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap B(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)+\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap B(X^{\prime},X)\ \Bigm{|}\ A(X,X^{\prime})\right)$
$=2\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\bigcap B(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)\;\text{($A(X,X^{\prime})$ and $C(X,X^{\prime})$ symmetric in $X$ and $X^{\prime}$)}$
$=2\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)\mathbb{P}_{X,X^{\prime}}\left(B(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)\ .$
We note that $\left(\ B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)+\left(\ \overline{B(X,X^{\prime})}\Bigm{|}A(X,X^{\prime})\right)=1$. Using Assumption 1, we have that $\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)+\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1$. As $A(X,X^{\prime})$ is symmetric in $X$ and $X^{\prime}$, we obtain that $2\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1$. Hence, $\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1/2$. Therefore,
$\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)$
$=2\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)\cdot\frac{1}{2}$
$=\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon\ \Bigm{|}\ A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)$
$=\frac{\mathbb{P}_{X,X^{\prime}}\left(\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon\bigcap A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}{\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}$
$\leq\frac{\mathbb{P}_{X,X^{\prime}}\left(\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon\right)}{\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}$
$=\frac{\mathbb{P}_{X}\left(\left|F_{1}(x,X)-F_{2}(x,X)\right|\geq\varepsilon\right)}{\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}\;\text{(numerator independent of $X^{\prime}$)}$
[…]
$\leq\frac{(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}\ .$
Note that by using Assumption 1 once again, we have
$\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)$ $=\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)+\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap\overline{B}(X,X^{\prime})\right)$
$=\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)+\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X^{\prime},X)\right)$
$=2\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)\ .$
Hence, we obtain
$\mathbb{P}_{X,X^{\prime}}\left(C(X,X^{\prime})\ \Bigm{|}\ A(X,X^{\prime})\right)\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}\ .$
This completes the proof. ∎
Appendix C Proof of the theorem on global coherence
C.1 Proof of Theorem 3.3
Theorem C.1 (Theorem 3.3 in the main text).
Assume that $f_{2}$ is globally $\alpha$-perception coherent with $f_{1}$ (i.e., $\mathbb{E}_{X}[\phi_{f_{1},f_{2}}(X)]\geq\alpha$). Let $X_{1}$, $X_{2}$ and $X_{3}$ be i.i.d. random variables following the law $\mathcal{D}_{\mathcal{X}}$. Then,
  1. 1.
    Global rank preservation. For all $\varepsilon_{2}>\varepsilon_{1}>0$, we have
    $\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\right|\geq\varepsilon_{2}\Bigm{|}\left|F_{1}(X_{1},X_{2})-F_{1}(X_{1},X_{3})\right|\leq\varepsilon_{1}\right)$
    $\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$
    where $C_{\varepsilon_{1}}$ is a positive constant depending on $\varepsilon_{1}$.
  2. 2.
    Global relative order preservation. For all $\varepsilon>0$, we have
    $\mathbb{P}_{X_{1},X_{2},X_{3}}\left(F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\geq\varepsilon\Bigm{|}F_{1}(X_{1},X_{2})\leq F_{1}(X_{1},X_{3})\right)\leq\frac{C(1-\alpha)}{\varepsilon}\ ,$
    where $C$ is a positive constant.
Proof.
1. Global rank preservation.
Let $A(X_{1},X_{2},X_{3})$ be the event $\left|F_{1}(X_{1},X_{2})-F_{1}(X_{1},X_{3})\right|\leq\varepsilon_{1}$. Let $\varepsilon=\frac{\varepsilon_{2}-\varepsilon_{1}}{2}$. Using the same technique as in the proof of Theorem 3.1, we can obtain that
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\right|\geq\varepsilon_{2}\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
$\leq\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\max\left(\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|,\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|\right)\geq\varepsilon\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
Applying Lemma 2, we have that
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\right|\geq\varepsilon_{2}\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}=\frac{4(1-\alpha)}{(\varepsilon_{2}-\varepsilon_{1})\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}$
$=\frac{4(1-\alpha)}{(\varepsilon_{2}-\varepsilon_{1})\mathbb{P}_{X_{1},X_{2},X_{3}}(\left|F_{1}(X_{1},X_{2})-F_{1}(X_{1},X_{3})\right|\leq\varepsilon_{1})}=\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$
where $C_{\varepsilon_{1}}=\frac{4}{\mathbb{P}_{X_{1},X_{2},X_{3}}(\left|F_{1}(X_{1},X_{2})-F_{1}(X_{1},X_{3})\right|\leq\varepsilon_{1})}$. This completes the proof.
2. Global relative order preservation. Let $A(X_{1},X_{2},X_{3})$ be the event $F_{1}(X_{1},X_{2})\leq F_{1}(X_{1},X_{3})$. Using the same technique as in the proof of Theorem 3.2, we can obtain that
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\geq\varepsilon\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
$\leq\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\max\left(\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|,\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|\right)\geq\frac{\varepsilon}{2}\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
Applying Lemma 2 leads to
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{2}(X_{1},X_{2})-F_{2}(X_{1},X_{3})\right|\geq\varepsilon\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
$\leq\frac{4(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}$
$=\frac{4(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}=\frac{C(1-\alpha)}{\varepsilon}\ ,$
where $C=\frac{4}{\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}$. This completes the proof. ∎
C.2 Lemma 2
Lemma 2.
Let $X_{1}$, $X_{2}$ and $X_{3}$ be the i.i.d random variables following the law $\mathcal{D}_{\mathcal{X}}$ and $A(X_{1},X_{2},X_{3})$ be an event depending on these three variables, $A$ is symmetric in $X_{2}$ and $X_{3}$. Assume that $f_{2}$ is $\alpha$-perception coherent with $f_{1}$ ($\mathbb{E}_{X}[\phi_{f_{1},f_{2}}(X)]\geq\alpha$). Then, for any $\varepsilon>0$, we have that
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\max\left(\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|,\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|\right)\geq\varepsilon\ \big{|}\ A(X_{1},X_{2},X_{3})\right)$
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\right)}\ .$ (C.1)
Proof.
For brevity, we shall use some compact notations. Let $B(X_{1},X_{2},X_{3})$ be the event $\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|\geq\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|$. We note that $B(X_{1},X_{3},X_{2})$ is the event $\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|\leq\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|$. Furthermore, let $C(X_{1},X_{2},X_{3})$ be the event $\max\left(\left|F_{2}(X_{1},X_{2})-F_{1}(X_{1},X_{2})\right|,\left|F_{1}(X_{1},X_{3})-F_{2}(X_{1},X_{3})\right|\right)\geq\varepsilon$. We have
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$=\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$+\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap\overline{B(X_{1},X_{2},X_{3})}\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)\ .$
Using Assumption 1, we have
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap\overline{B(X_{1},X_{2},X_{3})}\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$=\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{3},X_{2})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)\ .$
Therefore,
$\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$=\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$+\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{3},X_{2})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$=2\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\right)$
$=2\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\ \Bigm{|}\ A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)$
×$\times\mathbb{P}_{X_{1},X_{2},X_{3}}\left(B(X_{1},X_{2},X_{3})\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
$=2\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\Bigm{|}A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)\times 1/2$
$=\mathbb{P}_{X_{1},X_{2},X_{3}}\left(C(X_{1},X_{2},X_{3})\Bigm{|}A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)$
$=\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\Bigm{|}A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)$
$=\frac{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\bigcap A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}{\mathbb{P}_{X,X^{\prime}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$\leq\frac{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\right)}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$=\frac{\mathbb{P}_{X_{1},X_{2}}\left(\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\right)}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}\;\text{(numerator independent of $X_{3}$)}$
$=\frac{\mathbb{E}_{X_{1},X_{2}}\left[\mathds{1}_{\{\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\}}\right]}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$=\frac{\mathbb{E}_{X_{1}}\left[\mathbb{E}_{X_{2}}\left[\mathds{1}_{\{\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\}}\right]\right]}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$=\frac{\mathbb{E}_{X_{1}}\left[\mathbb{P}_{X_{2}}\left(\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\geq\varepsilon\right)\right]}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
[…]
$=\frac{\mathbb{E}_{X_{1}}\left[1-\phi_{f_{1},f_{2}}(X_{1})\right]}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$=\frac{1-\mathbb{E}_{X_{1}}\left[\phi_{f_{1},f_{2}}(X_{1})\right]}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$\leq\frac{1-\alpha}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}$
$=\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\right)}\ .$
This completes the proof. ∎
Appendix D Proof of Theorem 3.4
Theorem D.1 (Theorem 3.4 in the main text).
Consider a non-empty set $\mathcal{A}\subseteq\mathcal{X}$ such that $\phi_{f_{1},f_{2}}(x)\geq\alpha,\;\forall x\in\mathcal{A}$. For $0<\delta\leq 1$, let $\mathcal{A}_{\delta}=\{x\in\mathcal{X}:\min_{x^{\prime}\in\mathcal{A}}F(x^{\prime},x)\leq\delta\}$. Then, for any $\varepsilon$ ($0<\varepsilon\leq\alpha$), we have
$\mathbb{P}\left(\phi_{f_{1},f_{2}}(X)\leq\alpha-\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)\leq\frac{O(\delta)}{\varepsilon C_{\mathcal{A}_{\delta}}}\ ,$
where $C_{\mathcal{A}_{\delta}}$ is a positive constant depending on $\mathcal{A}_{\delta}$.
Proof.
Let us consider an $x\in\mathcal{A}_{\delta}$ and let $a(x)$ be an $\operatorname*{arg\,min}_{x^{\prime}\in\mathcal{A}}F(x^{\prime},x)$ (if not unique). We have that
$F_{1}(a(x),x)\leq\delta\;\text{and}\;F_{2}(a(x),x)\leq\delta\ ,$
which implies $d_{1}(x,a(x))=O(\delta)\;\text{and}\;d_{2}(x,a(x))=O(\delta)$. Notice that $O(\delta)$ is uniform on $x\in\mathcal{A}_{\delta}$ and $x^{\prime}\in\mathcal{X}$. Moreover, for any $x^{\prime}\in\mathcal{X}$, we have
$d_{1}(a(x),x^{\prime})-d_{1}(x,a(x))\leq d_{1}(x,x^{\prime})\leq d_{1}(a(x),x^{\prime})+d_{1}(x,a(x))\ .$
Hence,
$F_{1}(x,x^{\prime})$ $\leq\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(a(x),x^{\prime})+d_{1}(x,a(x))\right)$
$=\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(a(x),x^{\prime}))\right)+\mathbb{P}_{X}\left(d_{1}(a(x),x^{\prime})\leq d_{1}(x,X)\leq d_{1}(a(x),x^{\prime})+d_{1}(x,a(x))\right)$
$=F_{1}(a(x),x^{\prime})+O(d_{1}(x,a(x))=F_{1}(a(x),x^{\prime})+O(\delta)\ .$
On the other hand,
$F_{1}(x,x^{\prime})$ $\geq\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(a(x),x^{\prime})-d_{1}(x,a(x))\right)$
$=\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(a(x),x^{\prime}))\right)-\mathbb{P}_{X}\left(d_{1}(a(x),x^{\prime})-d_{1}(x,a(x))\leq d_{1}(x,X)\leq d_{1}(a(x),x^{\prime})\right)$
$=F_{1}(a(x),x^{\prime})-O(d_{1}(x,a(x))=F_{1}(a(x),x^{\prime})-O(\delta)\ .$
Therefore, $\left|F_{1}(x,x^{\prime})-F_{1}(a(x),x^{\prime})\right|=O(\delta)$, for any $x^{\prime}\in\mathcal{X}$. Proceeding in the same way, we also have $\left|F_{2}(x,x^{\prime})-F_{2}(a(x),x^{\prime})\right|=O(\delta)$. Thus,
$\left|F_{1}(x,x^{\prime})-F_{2}(x,x^{\prime})\right|$
$=\left|F_{1}(x,x^{\prime})-F_{1}(a(x),x^{\prime})+F_{2}(a(x),x^{\prime})-F_{2}(x,x^{\prime})+F_{1}(a(x),x^{\prime})-F_{2}(a(x),x^{\prime})\right|$
$\leq\left|F_{1}(x,x^{\prime})-F_{1}(a(x),x^{\prime})\right|+\left|F_{2}(a(x),x^{\prime})-F_{2}(x,x^{\prime})\right|+\left|F_{1}(a(x),x^{\prime})-F_{2}(a(x),x^{\prime})\right|$
$=\left|F_{1}(a(x),x^{\prime})-F_{2}(a(x),x^{\prime})\right|+O(\delta)\ .$
Therefore, there exists a positive constant $C_{\delta}=O(\delta)$ such that $\left|F_{1}(x,x^{\prime})-F_{2}(x,x^{\prime})\right|\leq\left|F_{1}(a(x),x^{\prime})-F_{2}(a(x),x^{\prime})\right|+C_{\delta}$, for any $(x,x^{\prime})\in\mathcal{A}_{\delta}\times\mathcal{X}$. Consequently,
$\mathbb{P}_{X}\left(\phi_{f_{1},f_{2}}(X)\leq\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)$
$=\mathbb{P}_{X}\left(1-\mathbb{E}_{X^{\prime}}\left[\left|F_{1}(X,X^{\prime})-F_{2}(X,X^{\prime})\right|\right]\leq\alpha-\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)$
$=\mathbb{P}_{X}\left(\mathbb{E}_{X^{\prime}}\left[\left|F_{1}(X,X^{\prime})-F_{2}(X,X^{\prime})\right|\right]\geq 1-\alpha+\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)$
$\leq\mathbb{P}_{X}\left(\mathbb{E}_{X^{\prime}}\left[\left|F_{1}(a(X),X^{\prime})-F_{2}(a(X),X^{\prime})\right|\right]+C_{\delta}\geq 1-\alpha+\varepsilon\ \Bigm{|}X\in\mathcal{A}_{\delta}\right)$
$=\frac{\mathbb{P}_{X}\left(\mathbb{E}_{X^{\prime}}\left[\left|F_{1}(a(X),X^{\prime})-F_{2}(a(X),X^{\prime})\right|\right]+C_{\delta}\geq 1-\alpha+\varepsilon\ \bigcap X\in\mathcal{A}_{\delta}\right)}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}$
$\leq\frac{\mathbb{P}_{X}\left(\mathbb{E}_{X^{\prime}}\left[\left|F_{1}(a(X),X^{\prime})-F_{2}(a(X),X^{\prime})\right|\right]+C_{\delta}\geq 1-\alpha+\varepsilon\right)}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}$
$=\frac{\mathbb{P}_{X}\left(1-\phi_{f_{1},f_{2}}(a(X))+C_{\delta}\geq 1-\alpha+\varepsilon\right)}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}$
$\leq\frac{\mathbb{P}_{X}\left(1-\alpha+C_{\delta}\geq 1-\alpha+\varepsilon\right)}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}\;(\phi_{f_{1},f_{2}}(a(X))\geq\alpha)$
$=\frac{\mathbb{P}_{X}\left(C_{\delta}\geq\varepsilon\right)}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}=\frac{\mathds{1}_{\{C_{\delta}\geq\varepsilon\}}}{\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}$
$\leq\frac{C_{\delta}}{\varepsilon\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}\;(\mathds{1}_{\{C_{\delta}\geq\varepsilon\}}\leq\frac{C_{\delta}}{\varepsilon})$
$=\frac{O(\delta)}{\varepsilon\mathbb{P}_{X}\left(X\in\mathcal{A}_{\delta}\right)}$
$\leq\frac{O(\delta)}{\varepsilon\mathbb{P}_{X}\left(X\in\mathcal{A}\right)}=\frac{O(\delta)}{\varepsilon C_{\mathcal{A}}}\ .$
This completes the proof of Theorem 3.4.
Appendix E Experiment details
For toy datasets on 2D and 3D, we simply use Euclidean distance as the dissimilarity metric. For all the experiments on neural networks, we use the cosine dissimilarity defined as
$d(u,v)=\frac{1}{2}\left(1-\frac{\langle u,v\rangle}{\|u\|\cdot\|v\|}\right)\ .$ (E.1)
All the experiments are conducted using a single GPU NVIDIA GeForce RTX 4090. The running time for each experiment is reported in each sub-section.
E.1 Toy dataset on 2D and 3D
For these experiments, we use scikit-learn package in Python to simulate the datasets. The two-moon dataset has 700 points, and the 5 $2D$ Gaussian clusters have 1000 points (Fig. E.1). The 3D clusters also have 1000 points. The student configuration is randomly initialized using Gaussian distribution. For training, we use batch size equal to 64, and learning rate equal to 0.1. We use Adam optimizer of the Pytorch package for training during 800 epochs. We set the temperature equal to $0.1$ in the soft ranking function.
Figure E.1: Matching process on $2D$ Gaussian datasets. We observe that learned configuration preserve a global structural coherence without preserving completely the geometry of the teacher configuration.
E.2 Retrieval experiments on CIFAR10 and CUB-200
Following [22], we use ResNet18 as the teacher model for both datasets. We also use the same student models for each datasets, shown in Fig. E.2. We apply our method only on the penultimate layer (dim 64 in ig. E.2), without using the final classification layer.
Training teacher model on CIFAR10. We use batch size equal to 64. We apply during training the data augmentation techniques, including random cropping (32x32, padding=4) and random horizontal flipping. We use the optimizer with following hyper-parameters: SGD(lr=0.1, momentum=0.9, nesterov=True, weight_decay=5e-4). We train the model from scratch for 100 epochs, where we decay the learning rate by a factor of 0.2 at epochs 40 and 80. The training time is less than 30 minutes.
Training teacher model on on the subset of CUB-200. We use batch size equal to 64. All images are resized to 256x256. We apply during training the data augmentation techniques, including random cropping (224x224), random horizontal flipping and random rotation (30 degrees). We use the optimizer with following hyper-parameters: SGD(lr=0.001, momentum=0.9, nesterov=True, weight_decay=5e-4). We train the model from the default pretrained model of Pytorch for 100 epochs, where we decay the learning rate by a factor of 0.1 at epoch 50. The training time is less than 5 minutes.
Transferring process. We use the training set as the transfer set, without using label. For both datasets, we train for 150 epochs. The technique is applied on the penultimate layer of teacher model (dimension of 512) and the penultimate layer of student models (dimension of 64). For temperature, we fix $\tau=0.1$ for teacher model and $\tau=0.3$ for student models. Notice that we use only the training set without label. We use the optimizer as follows: SGD(lr=0.05, momentum=0.9, nesterov=True, weight_decay=5e-4). We decay the learning rate by factor 0.1 at epochs 50 and 100. For each dataset, during transferring, we apply the same data augmentation techniques as for training the teacher model. The training time for CIFAR10 is about 20 minutes. The training time for the subset of CUB200 is less than 15 minutes.
(a) Student model for CIFAR10.
(b) Student model for CUB200.
Figure E.2: Student models for CIFAR10 and CUB200. Images are extracted from [22]. Notice that we perform transferring on the features of penultimate layer (dimension of 64), and do not use the final classification layer.
E.3 Classification on CIFAR-100
For teacher models, we use pretrained models released in the official code of the work [29] without any modification. We also use the same student models from the same code. We follow the same training scheme as the work [29]. More precisely, for optimizer, we use SGD(lr=0.01, momentum=0.9, weight_decay=5e-4). We use the training set for transferring, during 240 epochs, where we decay the learning by a factor of 0.1 at epochs 150, 180 and 210. The batch size is 64. Data augmentation techniques include random cropping (32x32, padding=4) and random horizontal flipping. We fix temperature $\tau=0.2$ for teacher model and $\tau=0.3$ for student models. Our technique is applied on the penultimate layer and the logit layer (before softmax function). The total loss for training the classification student model is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-entropy}}+\lambda\mathcal{L}_{\text{ours}}$, where we fix $\lambda=5$. We perform 5 independent runs to report results in Table 4.3. The running time for one run (240 epochs) is as follows: ShuffleNetV1: about 45 minutes. ShuffleNetV2: about 52 minutes. MobileNetV2: about 1 hour.
E.4 Ablation study: effect of batch size through qualitative evaluation
We use the toy datasets where we transfer from $3D\to 2D$ (Fig. E.3) and $2D\to 2D$ (Fig. E.4). We use different batch sizes ($B$) to observe the effect of batch size on the learned configuration. We see that when $B$ is too small ($B=3$), the learned configuration does not preserve a structural coherence of the reference one. This is to be expected, as a too small $B$ provides very little information about relation between data points. From $B=7$, the learned configuration preserves more or less the reference one. The results start to stabilize from $B=15$. It is reasonable to conjecture that we do not need too large $B$ (compared to the dataset size) to have the learned configuration preserving the coherence of the reference one. This is because in each training epoch, we randomly sample in each batch $B$ points, until finishing to process the whole dataset. Hence, each batch consider the relation between each point (reference point) to other $B-1$ points (compared set of points). Due to randomness of sampling, in each epoch, each point is compared to a different compared set of $B-1$ points. Hence, after different epochs, we should somehow have the global structural information between data points, without using a too large batch-size.
(a) Reference configuration.
(b) Batch size equal to 3.
(c) Batch size equal to 7.
(d) Batch size equal to 15.
(e) Batch size equal to 32.
(f) Batch size equal to 64.
(g) Batch size equal to 128.
Figure E.3: Ablation study of the effect of batch size on $3D\to 2D$ gaussian dataset, where we train with different batch sizes.
(a) Batch size equal to 3.
(b) Batch size equal to 7.
(c) Batch size equal to 15.
(d) Batch size equal to 32.
(e) Batch size equal to 64.
(f) Batch size equal to 128.
Figure E.4: Ablation study of the effect of batch size on two-moon dataset, where we train with different batch sizes.